The length of the top of the table is 6m greater that the width. The area is 72m^2. What is the width? What is the length?
Answer by checkley77(12844)

L=W+6
A=LW
72=(W+6)W
72=W^2+6W
W^2+6W-72=0
(W+12)(W-6)=0
W-6=0
W=6 ANS. FOR THE WIDTH.
L=6+6=12
PROOF:
72=12*6
72=72